# Towards Next Era of Multi-objective Optimization: Large Language Models as Architects of Evolutionary Operators

### v1

@@ IV-B4 Evaluation on MOTSPs
(b) MOTSP-3
(c) MOTSP-5
(d) MOTSP-6
(e) MOTSP-8
(f) MOTSP-9
@@ V Conclusion
In this work, we have embarked on a journey to harness the capabilities of LLMs for the evolution of evolutionary operators, specifically targeting the nuanced domain of multi-objective optimization problems. Our proposed framework, LLMOPT, represents a paradigm shift in EA operator design, moving towards an autonomous methodology that minimizes the need for expert intervention. The ’Pilot Run & Repair’ mechanism embedded within LLMOPT has been instrumental in refining the evolution process of EA operators, enhancing the robustness of the generated code and setting the stage for fully autonomous programming. The empirical studies showcases the superiority of LLM-evolved operators, demonstrating their advantage over traditional human-crafted approaches.
The future of this field is promising, with the potential for LLMs to further revolutionize the optimization landscape. Our results point to an emerging era of self-evolving, self-adapting, and self-improving programs, autonomously driven by the very challenges they are engineered to conquer. The application of our framework to a broader spectrum of MOPs, particularly those of greater complexity and scale, stands as a significant next step.

### v2

@@ IV-B3 Evaluation on MOTSPs
(b) MOTSP-5
(c) MOTSP-6
(d) MOTSP-9
@@
IV-B4 Efficiency of Algorithms
To evaluate the search efficiency of the EA operator generated via our proposed method, we focus on the execution time (wall clock time) as a primary metric, which is measured in seconds for a fair comparison across the algorithms MOEAD, CTAEA, RVEA, NSGA2, AGEMOEA, SMSEMOA. The investigation was conducted using testing groups of CMOP, MOKP, and MOTSP.
TABLE V: Running time (wall clock time measured in seconds) obtained by MOEAD, CTAEA, RVEA, NSGA2, AGEMOEA, SMSEMOA, and the best operator optimized through LLMOPT on CMOP, MOKP and MOTSP. Smaller running time on each category is highlighted in bold font.
CMOP MOKP MOTSP
LLMOPT 32 21 373
MOEAD 202 135
CTAEA 114 79 94
RVEA 37 54 64
NSGA2 23 45 53
AGEMOEA 50 52 60
SMSEMOA 1491 48 56
The empirical evaluation of the search efficiency, as presented in Table V, indicates a significant variance in execution time across the competing algorithms. While the best operator generated via LLMOPT excels in MOKP, achieving the lowest running time of 21 seconds, it falls behind in MOTSP, with a longer running time of 373 seconds, indicating the worst efficiency among the compared multi-objective algorithms. Despite this, the operator’s efficacy in terms of solution quality for MOTSPs is superior, suggesting that its strategic approach, while time-intensive, results in high-quality solutions. In the domain of CMOPs, the operator’s execution time is commendably efficient at 32 seconds, narrowly trailing behind NSGA2’s 23 seconds. This demonstrates the robustness of the proposed LLMOPT when handling different MOP categories.
IV-C Investigation on Different LLMs
In this section, we rigorously validate the performance of our proposed LLMOPT framework using various LLMs. For brevity, we denote them as GPT4, GPT4O, GEMINI, and CLAUDE, respectively. Moreover, the same temperature (0.5) is adopted for each LLM. Fig. 10 has illustrated the convergence curves obtained on CMOP validation dataset using these LLMs. As can be observed, different LLMs exhibit distinct behaviors within the proposed LLMOPT paradigm. Particularly, GPT4 and GPT4O initially struggle to produce a competent EA operator at the start of the search. However, through iterative optimization using LLMOPT, the effectiveness of operators generated by GPT4 and GPT4O improves significantly, highlighting the advantages of LLMOPT over traditional one-off program generation approaches. In contrast, CLAUDE consistently achieves the best EA operator performance from the outset until the end of the search. While GEMINI does yield a qualified EA operator initially, it fails to further enhance operator effectiveness over subsequent iterations. Based on these observations, the CLAUDE may have a superior performance in generating innovative EA operators against other LLMs.
Figure 10: Convergence curves gained by the proposed LLMOPT using different LLMs.
IV-D Ablation Study
In this section, the ablation studies are conducted to demonstrate the efficacy of the proposed Pilot Run & Repair component and dynamic selection strategy for LLM-based crossover.
IV-D1 In-depth Analysis of Code Quality Among LLMs
As we discussed in the introduction, a common challenge of using LLMs in program design is the occurrence of execution errors in the generated code, which can halt the iterative refinement process. To validate the efficacy of the proposed Pilot Run & Repair component, we tested the code quality generated by GPT4, GPT4O, GEMINI and CLAUDE, using the initialization prompts. Particularly, each LLM was tasked to generate 20 EA operators, which were then assessed for their operational viability. During the testing phase, any errors encountered triggered a repair protocol, which is designed to systematically address and resolve coding errors. The results have been summarized in Table VI, which presents a comparative analysis of the initial operator generation and subsequent repair success rates. As can be observed from, the ‘Tests’ column records the number of attempts made to generate functional EA operators. The “A + B” within the ‘Success’ column represents the total number of successfully generated EA operators, where “A” quantifies the instances without further interaction while “B” counts the successfully repaired instances via the proposed method. Moreover, ‘Failed’ tallies the number of unrepaired operators after one trial, while ‘Repaired’ reflects the proportion of successfully corrected operators through the repair process. The data presented in Table VI clearly demonstrates the superior initial success rates of GPT4O and GEMINI, with an impressive 80% of their generated operators being immediately deployable. Conversely, GPT4 shows a lower initial success but an outstanding repair success rate of 81.8%, underscoring the ‘Pilot Run & Repair’ component’s capability to effectively rectify errors. CLAUDE presents a balanced profile, showcasing a consistent performance in both initial operator generation and subsequent repairability.
TABLE VI: Quality of operator generated with different LLMs
LLM Tests Success Failed Repaired (%)
GPT4 20 9 + 9 2 81.8%
GPT4O 20 16 + 2 2 50.0%
GEMINI 20 16 + 3 1 75.0%
CLAUDE 20 10 + 7 3 70.0%
Figure 11: Convergence curves gained by the proposed LLMOPT (denoted by ‘GPT4’) and its counterpart without dynamic selection strategy (denoted by GPT4-no-dc).
IV-D2 Investigation on Dynamic Selection Strategy
The selectiopn strategy plays an important role in evolutionary progresses. To validate the performance of the developed dynamic selection strategy, tailored for LLM-based textual optimization tasks, we present the convergence curves of the proposed LLMOPT (denoted by ‘GPT4’) and its counterpart lacking this strategy (denoted by ‘GPT4-no-dc’). As can be observed in Fig. 11, the ‘GPT4’ model, equipped with the dynamic selection strategy, shows a significant improvement in terms of scores gained on CMOPs, indicating an effective adaptation and optimization capability brought by the developed dynamic selection strategy. In contrast, despite a good initialization for the operators, the ‘GPT4-no-dc’ model’s performance remains stagnant, again underscoring the necessity and impact of a dynamic selection strategy in the proposed LLMOPT.
@@ V Conclusion
In this work, we have embarked on a journey to harness the capabilities of LLMs for the evolution of evolutionary operators, specifically targeting the nuanced domain of multi-objective optimization problems. Our proposed framework, LLMOPT, represents a paradigm shift in EA operator design, moving towards an autonomous methodology that minimizes the need for expert intervention. The ’Pilot Run & Repair’ mechanism embedded within LLMOPT has been instrumental in refining the evolution process of EA operators, enhancing the robustness of the generated operator and setting the stage for fully autonomous programming. The empirical studies showcases the superiority of LLM-evolved operators, demonstrating their advantage over traditional human-crafted approaches.
The future of this field is promising, with the potential for LLMs to further revolutionize the optimization landscape. Our results point to an emerging era of self-evolving, self-adapting, and self-improving programs, autonomously driven by the challenges they are engineered to conquer. The application of our framework to a broader spectrum of MOPs, particularly those of greater complexity and scale, stands as a significant next step.